OpenStudy (anonymous):

The Baldwin family walked 4/5 of a mile in 4/9 of an hour. What is their unit rate in miles per hour?

OpenStudy (xhawaiiangirlx):

9/5 miles per hour is the answer

OpenStudy (anonymous):

Convert the statement into hours by multiplying the whole thing by 9/4.

OpenStudy (xhawaiiangirlx):

Wait (4/5) / (4/9) = 4/5 * 9/4 = 36/20 = 1 16/20 = 1 4/5 miles per hr